OpenStudy (anonymous):

Simplify by factoring √x9 (this x to the power 9)

OpenStudy (anonymous):

\[\LARGE \sqrt{x^9} =\sqrt{x^6\cdot x^3}=x^2\sqrt{x^3}\]

OpenStudy (anonymous):

is this what you're asking for?

OpenStudy (anonymous):

yes thanx --

OpenStudy (anonymous):

How do I use the key board to write x2√x3